OpenStudy (anonymous):

Which algebraic expression is equivalent to the expression below? 9(4x - 6)

OpenStudy (anonymous):

@heavenfun13

OpenStudy (anonymous):

@asnaseer

OpenStudy (asnaseer):

you need to distribute the 9 into the expression in brackets. so:\[9(4x-6)=9*4x-9*6=?\]